Find the binding energy of an e- in the grnd state of a H like atom in whose spectrum the third balmer line is equal to 108.5nm

Grade:12

1 Answers

vikas askiitian expert
13 years ago

1/lamda = R(1/22 - 1/52)Z2

 Z = 2

now energy of n=1 is = -13.6Z2 ev

binding energy = 13.6Z2 =13.6*4=54.4ev